# LaF-GRPO: In-Situ Navigation Instruction Generation for the Visually Impaired via GRPO with LLM-as-Follower Reward

### v2

@@ 1 Introduction
Figure 1: A NIG4VI data sample. (a) Bird’s-eye view map with waypoints and corresponding front-view images. (b) A comparison of instructions generated by ASSISTER, the ground truth, and our model (LaF-GRPO).
The Visually Impaired (VI) community, comprising approximately 2.2 billion (WHO2019vision) individuals globally with partial or complete blindness, underscores the significant need for effective assistive technologies. Extensive research in Visually Impaired Assistance (VIA) has emerged to address this need (zhao2024vialmsurveybenchmarkvisually; yuan2025walkvlmaidvisuallyimpairedpeople; gao2025wearable; cai2024navigating). This paper targets a critical and foundational sub-area: Navigation Instruction Generation for the Visually Impaired (NIG-VI). Unlike Navigation Instruction Generation (NIG) for general embodied agents which produces high-level trajectory plans (DBLP:conf/naacl/DouP22; DBLP:conf/acl/GopinathanMAS24; DBLP:conf/eccv/FanLWY24; DBLP:conf/eccv/KongCWSHYL24), NIG-VI is fundamentally human-centered. As illustrated in Figure 1, it demands the generation of in-situ, step-level instructions that must: (1) integrate non-visual sensory cues like auditory/tactile landmarks and surface textures; (2) provide precise directional and distance guidance to compensate for the lack of visual referencing; and (3) dynamically adapt to urban obstacles, all while ensuring user safety within a map-coordinate system.
@@
To bridge this gap, we propose LLM-as-Follower GRPO (LaF-GRPO), the first GRPO-based framework for the NIG-VI task. It features two novel components: (1) an LLM that simulates how VI users respond to navigation instructions by interpreting their likely actions, and (2) a VLM post-training procedure for instruction generation guided by an LLM-as-Follower reward. LaF-GRPO mitigates the need for costly trials with VI users while ensuring instruction usability through a human-in-the-loop navigation simulation. Furthermore, to address the scarcity of VI navigation benchmarks, we introduce NIG4VI—a comprehensive VI navigation instruction benchmark featuring 27k samples for simulation experiments. Fully open-sourced and annotated with granular spatial metadata, NIG4VI enables the generation of detailed, open-ended, in-situ instructions.
Evaluation of LaF-GRPO on the NIG4VI benchmark reveals the following findings: (1) Qwen2.5-VL models under the Zero-(LaF-GRPO) paradigm significantly outperform the standard zero-shot baseline, demonstrating superior performance across a range of metrics. (2) Qwen2.5-VL-7B models trained with SFT+(LaF-GRPO) show leading performance, achieving a METEOR score of 0.542, substantially surpassing powerful proprietary models like GPT-4o. (3) Beyond quantitative improvements, qualitative analysis further confirms that LaF-GRPO generates more human-centered instructions, distinguished by enhanced linguistic variety, more intuitive directional cues, richer environmental details, and crucial safety considerations. In summary, our main contributions are:
@@ 6 Results and Discussions
Figure 4: A comparative case study of navigational guidance provided by SFT and SFT+(LaF-GRPO) methods across successive steps. Findings indicate that SFT+(LaF-GRPO) (Ours) generates instructions with greater linguistic variety and more effectively incorporates o’clock directions and specific travel distances.

### v3

@@ 1 Introduction
Figure 1: NIG4VI sample. (a) Bird’s-eye map and front views. (b) Instructions generated by ASSISTER (DBLP:conf/eccv/HuangSZBBO22), the ground truth, and LaF-GRPO (Ours).
The Visually Impaired (VI) community, comprising approximately 2.2 billion (WHO2019vision) individuals globally with partial or complete blindness, underscores the significant need for effective assistive technologies. Extensive research in Visually Impaired Assistance (VIA) has emerged to address this need (zhao2024vialmsurveybenchmarkvisually; yuan2025walkvlmaidvisuallyimpairedpeople; gao2025wearable; cai2024navigating). This paper targets a critical and foundational sub-area: Navigation Instruction Generation for the Visually Impaired (NIG-VI). Unlike Navigation Instruction Generation (NIG) for general embodied agents which produces high-level trajectory plans (DBLP:conf/naacl/DouP22; DBLP:conf/acl/GopinathanMAS24; DBLP:conf/eccv/FanLWY24; DBLP:conf/eccv/KongCWSHYL24), NIG-VI is fundamentally human-centered. As illustrated in Figure 1, the task demands in-situ, step-level instructions that (1) incorporate non-visual cues, (2) ensure precise directional and distance guidance, and (3) adapt to obstacles for safety in map coordinates.
@@
To bridge this gap, we propose LLM-as-Follower GRPO (LaF-GRPO), a novel GRPO-based framework for the NIG-VI task. It features: (1) an LLM that simulates how VI users respond to navigation instructions by interpreting their likely actions, and (2) a VLM post-training procedure for instruction generation guided by an LLM-as-Follower reward. LaF-GRPO mitigates the need for costly trials with VI users while ensuring instruction usability through a human-in-the-loop navigation simulation. Furthermore, to address the scarcity of VI navigation benchmarks, we introduce NIG4VI—a comprehensive VI navigation instruction benchmark featuring 27k samples. Fully open-sourced and annotated with granular spatial metadata, NIG4VI enables the generation of detailed, open-ended, in-situ instructions.
Evaluation of LaF-GRPO on the NIG4VI benchmark yields three key findings: (1) Qwen2.5-VL models under the Zero-(LaF-GRPO) paradigm significantly outperform the standard zero-shot baseline across multiple metrics. (2) Qwen2.5-VL-7B models trained with SFT+(LaF-GRPO) achieve state-of-the-art performance, reaching a METEOR score of 0.542 and substantially surpassing strong proprietary models such as GPT-4o. (3) Beyond quantitative gains, qualitative analysis shows that LaF-GRPO produces more human-centered instructions, characterized by greater linguistic diversity, more intuitive directional cues, richer environmental details, and essential safety considerations. In summary, our main contributions are:
@@ 6 Results and Discussions
Figure 4: A comparative case study of navigational guidance provided by SFT and SFT+(LaF-GRPO) methods.